a line contains the points (4,2) and 0,1). what is the equation of the line?



Answer by Cromlix(4381)

Hi there,
Gradient (Slope) = y2 - y1/x2 - x1
Gradient (Slope) = 1 - 2/0 - 4 = -1/-4 = 1/4
Using line equation y - b = m(x - a)
m = 1/4
(a,b) = (4,2)
y - 2 = 1/4(x - 4)
y - 2 = 1/4x - 1
y = 1/4x - 1 + 2
y = 1/4x + 1
or multiply through by 4
4y = x + 4
Hope this helps :-)